OpenStudy (algtrigcalc):

|2(1+k)-3|=|0.4k-3| How to find the solution of an absolute value equation?

OpenStudy (muratovic):

You can have two possible outcomes, 1) left side is equal to the right side (without absolute values) 2) left side is equal to the negative of the right side (without absolute values)

OpenStudy (algtrigcalc):

Do the absolute values of both sides cancel out? 2(1+k)-3=0.4k-3

OpenStudy (sshayer):

\[\pm(2(1+k)-3)=\pm (0.4k-3)\] either 2(1+k)-3=0.4k-3 (1) or 2(1+k)-3=-(0.4k-3) (2) or -{2(1+k)-3}=0.4k-3 i.e.,2(1+k)-3=-(0.4k-3) which is same as (2) or -{2(1+k)-3}=-(0.4k-3) i.e.,291+k)-3=0.4k-3 which is same as (1)

OpenStudy (sshayer):

correction i.e.,2(1+k)-3=0.4k-3

OpenStudy (sshayer):

solve only (1) and (2) for k

OpenStudy (algtrigcalc):

Oh I see. Thanks!!